OpenStudy (jh99):

explain 2cos^2(4x)-1=0

ganeshie8 (ganeshie8):

\(\cos(8x) = 0\) \(8x = \pi/2 + 2k\pi, ~~~~~3\pi/2 + 2n\pi\)

OpenStudy (sshayer):

\[8x=\frac{ \pi }{ 2 }+n \pi\] where n is an integer.

OpenStudy (holsteremission):

Alternatively, \[2\cos^2(4x)-1=\left(\sqrt2\cos(4x)-1\right)\left(\sqrt2\cos(4x)+1\right)=0\]giving you two equations to solve, \[\begin{cases}\sqrt2\cos(4x)-1=0\\[1ex]\sqrt2\cos(4x)+1=0\end{cases}\]More work than necessary, but it doesn't hurt :P

OpenStudy (shamim):

U know Cos(2A)=2*cos^2A-1

OpenStudy (shamim):

So plz tell me Cos (8A)=?